Operations Report Apr 21st

Crew 245 Operations Report 21APR2021

SOL: 11

Name of person filing report: Shravan Hariharan

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Generator: Working nominally

Hours run: 7

From what time last night: 9:30 pm, SOC 60%

To what time this morning: 4:30 am, SOC 100%

List any additional daytime hours when the generator was run: N/A

Solar— SOC 65% at 7:48 pm (Before generator is run at night)

Notes on power system: For the last several sols, the freezers in the Science Dome have been running in order to preserve science samples. Due to this, the station power consumption has been significantly higher than it previously was, resulting in the generator being turned on earlier (as early as 7:00 pm on colder days, and 9:00 pm on warmer days) once the SOC reaches below 70%.

Diesel Reading – 50%

Station Propane Reading – 70%

Water (loft tank): 25 gallons

Water Meter: 150665.5 units

Water (static tank): 135 gallons

Static to Loft Pump used – yes

Water in GreenHab: 0 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Spirit rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: yes, in town

Opportunity rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: no, in town

Curiosity rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: yes, in town

Perseverance rover used: No

Hours: 204.6

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Left tail-light on Perseverance rover is not functioning – noticed on EVA.

ATV’s Used: 350.2

Reason for use: EVA

Oil Added? No

ATV Fuel Used: 0.5 gallon (estimated)

# Hours the ATVs were used today: 4 (but 0.5 hour of active transit)

Notes on ATVs: Mission Director will monitor and refuel ATVs for duration of mission.

HabCar used and why, where?: Nothing to report

CrewCar used and why, where?: Nothing to report

General notes and comments: Nothing to report

Summary of internet: Internet is working nominally. Crew is rationing internet usage during the day to ensure that there is sufficient bandwidth for Mission Support Comms window.

Summary of suits and radios: All radios nominal, all suits currently charging. Suits 7, 9, and 11 used on EVA today, and airflow reduced for all suits approximately 3 hours into EVA. No further EVAs planned with these suits, but Crew Engineer suspects issue is with the chosen charging outlet – all three suits were charging from same power strip. Crew Engineer will test these suits tosol to determine root cause of issue and any potential fixes.

Summary of Hab operations: Flapping side panel on outside of Hab. Station Director secured panel with 2×4, so it is no longer making noise

Summary of GreenHab operations: Being used to host Crew Botanist regolith feasibility experiment, but no active operations.

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

EVA Report – Apr 21st

Crew 245 EVA Report 21APR2021

EVA # 10

Author: Crew Scientist Olivia Ettlin

Purpose of EVA: Gathering soil samples and scouting for emergency shelters.

Start time: 1015 MDT

End time: 1330 MDT

Narrative: The EVA this morning started slightly after the beginning of the approved EVA window, as the crew was carefully gathering and sterilizing required scientific instruments according to XO Hume’s Planetary Protection Protocols. After we suited up and completed our 5-minute decompression in the airlock, we loaded our supplies into the Perseverance rover. Commander Dickstein and Health and Safety Officer Coultrup drove the Rover and Crew Scientist Ettlin took the ATV Serial Number 2. Starting charge on the Rover was 100% with 204.1 hours and the ATV required some choke before starting.

We traveled from the Hab to Entrance Road and proceeded south on Cow Dung Road for approximately 10 minutes until we reached Zubrin’s Head. Upon arrival at Zubrin’s head, we parked our vehicles on the side of the road. Using the analog map, we confirmed that our travel calculations had been accurate. From there, we embarked southwest on foot to explore the immediate region and investigate points of interest, which were communicated to us by our crewmates after their EVA to this site. As always, we carried the EVA med kit with us.

When we had completed gathering the samples, we returned to our vehicles and proceeded north on Cow Dung Road to Robert’s Rock Garden, where we did a final visit to drop waypoints and collect rock samples. From there, we returned to our vehicles and proceeded to Entrance Road, where we paused for approximately 20 minutes to capture some memorable photos with the MDRS entrance sign and facility as a backdrop, as directed by Mars Society Media Director Michael Stoltz. Finally, we returned back to the Hab. Upon returning, the Rover displayed 204.6 hours and 100% charge.

Destination: Zubrin’s Head

Coordinates of Destination: 12S 519500 UTM 4248100

Participants: Crew Scientist Olivia Ettlin, Commander Dylan Dickstein, and Health and Safety Officer Alex Coultrup

Road(s) and routes per MDRS Map: Hab to Entrance Rd, then South on Cow Dung road to Zubrin’s Head. Cow Dung Road to Entrance Rd, then back to the Hab.

Mode of travel: 2 team members in Perseverance Rover, 1 team member following close behind on ATV serial #2. All three members of the team walked the destination from the location where we parked the rovers safely on the side of the road.

Sol Summary Report Apr 21st

Crew 245 Sol Summary Report 21APR2021

Sol: 11

Summary Title: Approaching the Finish Line

Author’s name: Shayna Hume

Mission Status: Executive Officer

Sol Activity Summary:

· 08:30-09:30: Misc. breakfast and briefing

· 09:30-10:00: Prepare for EVA.

· 10:00-14:00: EVA.

· 14:00-15:00: Lunch! Jambalaya.

· 15:00-18:00: Separate science activities & rest time.

· 18:00-18:30: Crew Photos.

· 18:30-20:00: Dinner.

Look Ahead Plan: Tomorrow we have the final EVA of the mission – Crew #2.

Anomalies in work: None.

Weather: Nominal.

Crew Physical Status: No issues.

EVA: EVA #10 report coming in.

Reports to be filed: Operations Report, EVA Request, EVA Report

Support Requested: n/a

EVA Report – April 17th

Author: Julio A. Hernandez

Purpose of EVA: A training EVA to improve crew mapping abilities

Start time: 09:00

End time: 11:30

Narrative: The EVA this morning started exactly at the beginning of the approved EVA window. The crew took enhanced measures during the EVA debriefing to ensure that the EVA participants were well prepared for the mission. In short, the route was mapped and landmarks were identified to aid in navigation. The health and safety checks were conducted without any issue to note; the participants were engaged and eager for their EVAs. Crew Botanist Hernandez and Crew Engineer Hariharan were the driver and navigator in the Perseverance Rover, respectively. As per Outpost’s recommendation, ATV #2 was taken for the EVA by XO Hume. The starting charge of the Rover was 100% with 202.1 hours. The ATV started up after some difficulty due to a cold engine.

We proceeded toward our destination, Robert’s Rock Garden, following the Cow Dung Road, as shown on the map and per the pre-EVA briefing. The EVA team checked COMMs with HABCOM until they were out of range. The EVA team stopped on the north side of Robert’s Rock Garden and took a photo of Kissing Camel to the East. Phobos Peak was included in the photo.

After visiting Robert’s Rock Garden, the EVA team departed to Zubrin’s Head. Identifying Zubrin’s Head proved to be challenging. Using both the map, approximated GPS coordinates, and location of the Sun, the EVA team identified a feature that they believed to be Zubrin’s Head. As a precaution to ensure the identified geological feature was indeed Zubrin’s Head, the EVA team proceeded up and down Cow Dung Road to raise their confidence in their feature identification (the feature looks like a person wearing a hat?). The EVA team proceeded to take photos of themselves next to the structure.

The EVA team then proceeded to hike in a southwest direction to explore the area. Using a local small hill, the EVA team took photos of the Barrainca Butte. After this little excursion, the EVA team returned to their vehicles to relocate themselves to the south side of Robert’s Rock Garden to collect soil samples at four different locations before returning to HAB grounds. The EVA team was granted permission to re-enter the HAB ground by HABCOM. The vehicles were parked. The Rover was parked and 86% with 202.7 hours. The EVA team was granted permission to re-enter the HAB and the airlock was sealed and re-pressurized.

Destination: Robert’s Rock Garden, Robert’s Head, Barrainca Butte

Coordinates (use UTM NAD27 CONUS): Moving through 12 s 518250 utm 4249250 (Robert’s Rock Garden), parking at 12 s 519500 utm 4248100 (Robert’s Head), hiking to near 519000 4247000 (Barrainca Butte)

Participants: Shayna Hume, Shravan Hariharan, Julio Hernandez

Road(s) and routes per MDRS Map: South Cow Dung Road

Mode of travel: Vehicles and Walking

EVA # 7

Author: Health and Safety Officer Alex Coultrup

Purpose of EVA: Training and confirmation of orienteering skills and ability to navigate to specified destinations using analog maps.

Start time: 1230 MDT

End time: 1520 MDT

Narrative: The EVA this afternoon started right at the beginning of the approved EVA window. HSO performed our required health and safety checks (blood pressure, heart rate, and temperature checks) and carefully implemented our planetary protection measures as per XO Hume’s research protocol. After we suited up and completed our 5-minute decompression in the airlock, we loaded our supplies into the Perseverance rover. Commander Dickstein and Crew Scientist Ettlin drove the Rover and HSO Coultrup took the ATV Serial Number 2. Starting charge on the Rover was 89% with 202.8 hours and the ATV started up quickly with little choke needed.

We traveled from the Hab to Entrance Road, and pulled over to consult our map. By using the map to estimate the distance, and by estimating the speed of the rover, we were able to estimate the total expected time to travel between this location and our first destination. We resumed traveling north on Cow Dung Road for the duration we calculated, and arrived at Streambed Connector as anticipated. We parked the rovers safely on the side of the road. Using the same calculation techniques as before, we estimated the duration of our walk to Destination 1. Using the analog map, we walked to Destination 1, took a photo, and enjoyed the view of the beautiful rock strata to our north upon arrival. As always, we carried the EVA med kit with us for this walk. We returned back to the vehicles using the same route, and proceeded north on Cow Dung Road toward Destination 2.

While traveling toward Destination 2, we pulled over at Galileo Road to consult our map and ensure accurate navigation. After performing our time/distance calculations, we continued North along Cow Dung Road, and were able to visually determine when we passed Cowboy Corner. We continued slightly further down Cow Dung Road, safely parked the rovers, and walked a few hundred meters toward the reservoir area, where we captured a photo of our arrival. We returned back to the vehicles using the same route, and consulted our map to perform the same navigation process we had been employing so far. We determined the distance, speed, and rate of travel, and proceeded north on Cow Dung Road toward Destination 3.

As we approached toward Destination 3, we pulled at the end of our approximated travel duration, but by consulting the analog map and visual landmarks, we were able to determine that we were not yet at the desired location. We re-started our vehicles and proceeded 30 seconds further north, and reached the intersection of Tank Wash Road and Cow Dung Road. We took a photo facing west, and returned to our vehicles.

From Destination 3, we began travelling south on Cow Dung Road until we reached the Streambed Connector. At the Streambed Connector area, we pulled over once again and spent some time collecting soil samples for our scientific research. When we had completed gathering the samples, we returned to the vehicles and proceeded south on Cow Dung Road to Entrance Road, and back to the Hab. Upon returning, the Rover displayed 203.2 hours and 86% charge.

Destination 1: Beautiful strata north of canyon at end of Streambed Connector

Coordinates of Destination 1: 12 S 0519556, UTM 4250370

Destination 2: Reservoir Dam/waterhole

Coordinates of Destination 2: 12 S 0518579, UTM 4252676

Destination 3: Intersection of Tank Wash Road and Cow Dung Road

Coordinates of Destination 3: 12 S 0518522, UTM 4253399

Participants: Crew Scientist Olivia Ettlin, Commander Dylan Dickstein, and Health and Safety Officer Alex Coultrup

Road(s) and routes per MDRS Map: Hab to Entrance Rd, then North on Cow Dung road to Streambed Connector. Cow Dung Road to Galileo Road. Cow Dung Road to Cowboy Corner. Cow Dung Road to Tank Wash Road intersection. On return route took Cow Dung Rd South to Streambed Connector. Cow Dung Road to Entrance Rd, then back to the Hab.

Mode of travel: 2 team members in Perseverance Rover, 1 team member following close behind on ATV serial #2. All three members of the team walked to each destination from the locations where we parked the rovers safely on the side of the road.

Operations Report – April 17th

[category operations-report]

Crew 245 Operations Report 17-04-2021

SOL: 7

Name of person filing report: Shravan Hariharan

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Generator: Working nominally

Hours run: 7

From what time last night: 10:00 pm, SOC 57%

To what time this morning: 5:00 am, SOC 100%

List any additional daytime hours when the generator was run: N/A

Solar— SOC 45% at 7:40 pm (Before generator is run at night),

Diesel Reading – 50%

Station Propane Reading – 73%

Water (loft tank): 15 gallons

Water Meter: 150460.4 units

Water (static tank): 340 gallons

Static to Loft Pump used – yes

Water in GreenHab: 0 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: No

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Spirit rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: yes, in town

Opportunity rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: no, in town

Curiosity rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: yes, in town

Perseverance rover used: Yes

Hours: 203.3

Beginning charge: 100%

Ending charge: 86%

Currently charging: Yes

Notes on rovers: Nothing to report

ATV’s Used: 350.2

Reason for use: EVA

Oil Added? No

ATV Fuel Used: 1 gallon (estimated)

# Hours the ATVs were used today: 6 (but 1.5 hours of active transit)

Notes on ATVs: Mission Director will monitor and refuel ATVs for duration of mission

HabCar used and why, where?: No

CrewCar used and why, where?: No

General notes and comments: Nothing to report

Summary of internet: Internet is working nominally. Crew did not consume allotted bandwidth last night so there was an additional bandwidth bonus available today.

Summary of suits and radios: All radios nominal, all suits currently charging. Suits 3, 4, 6, 7, 9 and 11 all used for EVA today. Suit 7 battery would not charge after EVA, and Suit 9 battery was not producing sufficient power output even when fully charged, so both were replaced.

Summary of Hab operations: Nothing to report besides nominal crew activities.

Summary of GreenHab operations: Being used to host Crew Botanist botany experiment. Crew Botanist briefly checked plant growth and health in GreenHab today.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report.

Sol Summary – April 17th

Crew 245 Sol Summary Report 17-04-2021

Sol: 7

Summary Title: Twin EVA’s

Author’s name: Shayna Hume

Mission Status: Executive Officer

Sol Activity Summary:

· 07:30-08:30: Misc. breakfast with music while getting ready for EVA’s.

· 08:30-09:00: EVA Checklist.

· 09:00: Finish 5 minutes in airlock and EVA Crew #2 began EVA.

· 09:05-11:30: EVA went as planned.

· 11:30-12:00: Crew debrief.

· 12:30-12:35: Finish 5 minutes in airlock and EVA Crew #1 began EVA.

· 12:35-15:30: EVA went as planned.

· 15:30-17:30: Small science study checks and report writing.

· 17:30-18:30: Making lentil curry for dinner!

Look Ahead Plan: Tomorrow no EVA’s, but instead a focus on science and life in the hab!

Anomalies in work: None.

Weather: Nominal.

Crew Physical Status: No issues.

EVA: EVA #6-7. Reports incoming.

Reports to be filed: Operations Report, EVA Report

Support Requested: n/a

Shayna Hume, Executive Officer
Red Planet People – MDRS Crew 245 "Team Patamars"To Mars and Beyond – For All!
Instagram, Twitter, Facebook

uc?id=1aRg1AKKbJAbceXaUkRwv91ac9H3hhnnz&export=download